find two consecutive odd integers such that 3 times the smaller integer increased by 6 is equal to 4 times the larger integer decreased by 20
Answer by ankor@dixie-net.com(22740)

"find two consecutive odd integers"
x, (x+2)
"such that 3 times the smaller integer increased by 6 is equal to 4 times the larger integer decreased by 20"
3x + 6 = 4(x+2) - 20
:
3x + 6 = 4x + 8 - 20
:
3x + 6 = 4x - 12
:
6 + 12 = 4x - 3x
:
x = 18, & 20; which are an even integers, there are no odd integer solutions
:
:
Check solution in the statement:
"3 times the smaller integer increased by 6 is equal to 4 times the larger integer decreased by 20"
3(18)+ 6 = 4(20) - 20
54 + 6 + 80 - 20